OpenStudy (anonymous):

Solve for y dy/dx=sqrt(yx),y(1)=4 been waiting for about an hour for 1 question. is a little help too much to ask for? all the easy questions get answered and the hard ones are ignored because people are too damn lazy. the chat is completely offtopic the whole time ive been here and seems like new members will be ignored for got knows how long before anything is answered

OpenStudy (cruffo):

Maybe you can help me a little. I'm confused with the problem: \[\frac{dy}{dx}=\sqrt{yx}\] and you are given that \(y(1)=4\). What are you trying to solve for?

OpenStudy (anonymous):

solve for y equation should be in y = f(x) + C ive gotten to integral (dy/sqrt(y)) = integral(sqrt(x)dx) but after that all my answers have been wrong

OpenStudy (anonymous):

its the initial value problems

zepdrix (zepdrix):

 Hmm it sounds like you're on the right track.